Upton Chuck is riding the Giant Drop at Great America. If Upton free falls for 2.60 seconds, what will be his final velocity and how far will he fall?
1
Expert's answer
2020-05-20T09:09:47-0400
"v=gt=(9.8)(2.6)=25.5\frac{m}{s^2}"

"h=0.5gt^2=0.5(9.8)(2.6)^2=33.1\ m"